# On admissibility in post-hoc hypothesis testing

### v2

@@ Abstract
The validity of classical hypothesis testing requires the significance level $\alpha$ be fixed before any statistical analysis takes place. This is a stringent requirement. For instance, it prohibits updating $\alpha$ during (or after) an experiment due to changing concern about the cost of false positives, or to reflect unexpectedly strong evidence against the null. Perhaps most disturbingly, witnessing a p-value $p\ll\alpha$ vs $p=\alpha-\epsilon$ for tiny $\epsilon>0$ has no (statistical) relevance for any downstream decision-making. Following recent work of Grünwald (2024), we develop a theory of post-hoc hypothesis testing, enabling $\alpha$ to be chosen after seeing and analyzing the data. To study “good” post-hoc tests we introduce $\Gamma$-admissibility, where $\Gamma$ is a set of adversaries which map the data to a significance level. We classify the set of $\Gamma$-admissible rules for various sets $\Gamma$, showing they must be based on e-values, and recover the Neyman-Pearson lemma when $\Gamma$ is the constant map. We also give a Rao-Blackwellization result, proving that the expected utility of an e-value can be improved (for any concave utility) by conditioning on a sufficient statistic.
@@ 1 Introduction
In other words, the current paradigm of hypothesis testing requires that the significance level $\alpha$ be chosen independently of (thus without looking at) the data. This is well-known to statisticians, but continues to be a source of frustration and confusion for practitioners, and the epidemiologist’s mistake above is unfortunately all too common (Bakan, 1966; Dar et al., 1994; Gigerenzer, 2004). This fact is limiting: once the data has been gathered, if the selected value of $\alpha$ gives a vacuous or uninteresting result, nothing more can be done. The same data should not be used in any further analysis of the same question (or, technically, even a separate question motivated by the first [^1] This—admittedly philosophically thorny and somewhat controversial—reliance of p-values on counterfactuals is one of their many restrictive features (Pratt, 1964; Wagenmakers, 2007).).
The difficulties do not end there. Requiring that $\alpha$ be chosen independently of the data introduces a fundamental tension between evidence and decision-making. Namely, upon observing a p-value $p\leq\alpha$, it is irrelevant whether $p$ is roughly the same as $\alpha$ or significantly smaller. For example, if $\alpha=0.05$, then observing $p=0.04$ and $p=10^{-6}$ have the same decision-theoretic consequences. As pointed out recently by Grünwald (2024), this is a frustrating fact: the p-value is ostensibly a measure of evidence against the null hypothesis, yet any $p$-value below $\alpha$ has no formal relevance for any downstream decision task.
This limitation serves as an uncomfortable reminder that current statistical practice consists of an odd combination of both Ronald Fisher’s perspective and that of Jerzy Neyman and Egon Pearson. Whereas Fisher proposed the p-value as measure of evidence (Fisher, 1925, 1935), Neyman and Pearson were focused on decision making (Neyman and Pearson, 1928). The fact that we both report the p-value and make a decision at a prespecified significance level is an awkward methodological hybrid that none of these pioneers promoted.
This hybrid approach reflects (or perhaps introduces) confusion over the proper interpretation of p-values, significance levels, and hypothesis testing in general. And such confusion has substantial practical consequences. “The problem of roving alphas” as Goodman (1993) puts it, i.e., choosing $\alpha$ as a function of the data, is well-documented in business and management science (Hubbard and Bayarri, 2003; Hubbard, 2011), medicine (Goodman, 1999), and psychology (Gigerenzer, 2014). What are statisticians to do in the face of this mistake?
@@
Faced with such examples, in this work we consider an alternative response to the problem of roving alphas. Instead of pointing out flaws in the epidemiologist’s approach, we focus on building new statistical tools which function both as measures of evidence and allow for mathematically sound downstream decision-making. In particular, we develop the theory of post-hoc hypothesis testing (Grünwald, 2024), which allows for data-dependent significance levels. The catch is that the guarantees that an analyst can give on their procedure changes under this theory, moving from error probabilities to expected losses. If $\alpha$ is fixed beforehand, however, then the theory of post-hoc testing recovers the standard theory of hypothesis testing. Thus, in that sense, post-hoc hypothesis testing is a strict generation of the classical framework. Our focus in this paper is to define and classify “admissible test families”, which are the analogue of uniformly most powerful tests in the post-hoc setting.
One can situate our work as belonging to a recent push to resolve several issues with traditional statistical tools. The inability to handle post-hoc significance levels is just one problem with much of modern statistics—others include the inability to handle optional stopping and optional continuation. The burgeoning area of sequential, anytime-valid inference (SAVI) (Ramdas et al., 2023) is focused on such issues. A fundamental tool in this line of work is the e-value (Grünwald et al., 2024; Ramdas and Wang, 2025), an object which will also play a crucial role in this paper. Overall, we view our work here as an extension of SAVI: broadly, as an attempt to refine existing statistical technology to make life easier for practitioners and to mitigate the risk of statistical malpractice.
@@ 1.1 Contributions
Following Grünwald et al. (2024), in order to formulate a theory of post-hoc hypothesis testing we begin by following Wald (1939) and recast the usual notions of type-I and type-II error probabilities in terms of expectations over data-dependent type-I and type-II loss functions. We require that a test has bounded type-I loss in expectation, where the type-I loss may be chosen adversarially. Such a test is called type-I risk safe, a notion which generalizes the property of a test having a bounded type-I error probability.
At this point we depart from the setting studied by Grünwald (2024) and introduce a different notion of admissibility. We summarize our contributions below. While the general setting and several of the results are presented for composite hypotheses, we are mainly focused on testing point nulls and alternatives.
Throughout the paper it will often be useful to remark on the differences or similarities between our post-hoc setting and the traditional, non-post-hoc setting. Henceforth, we will often refer to the latter as the “traditional setting” or sometimes as the “Neyman-Pearson paradigm.” We refer to Neyman and Pearson because it was they who introduced type-I and type-II errors and the notion of a uniformly most powerful test (Neyman and Pearson, 1928, 1933), concepts that we build upon and generalize in this work.
@@
Definition 2.4 introduces the notion of $\Gamma$-admissibility, where $\Gamma$ is a set of functions which map the data to a type-II loss function. This generalizes the definition of a uniformly most powerful test to the post-hoc setting (Observation 2.5 makes this precise). Informally, a test is $\Gamma$-admissible if there is no test that has smaller expected type-II loss under each distribution in the alternative and across all mappings in $\Gamma$. Our definition of $\Gamma$-admissibility departs from the notion of admissibility studied by Grünwald (2024), which requires that a test be dominated with probability 1 under each distribution in the alternative to be inadmissible. We require that a test be dominated only in expectation, making fewer tests admissible.
@@
We give an extension of the likelihood ratio test—shown to be uniformly most powerful in the traditional paradigm (Neyman and Pearson, 1933)—to a post-hoc setting and prove that the extension is $\Gamma$-admissible for any $\Gamma$. This is in Section 3. This shows that our definition of admissibility is not too strict; the set of $\Gamma$-admissible tests is non-empty. This investigation also sheds light on an interesting phenomenon in the post-hoc setting: Even for continuous distributions, if randomization is allowed then admissible tests will always be randomized. More specifically, there will be losses on which the test is randomized. Since randomized tests may not be kosher for all applications, this motivates explicitly defining and studying the admissibility of binary post-hoc tests which reject or accept with probability one.
@@
Moving beyond the likelihood ratio test, we prove several general properties for all $\Gamma$-admissible tests for any $\Gamma$. We study both randomized and binary tests. Of particular note, we associate with each test an e-variable and show how any admissible test can be written as a function of the associated e-variable. Thus, the e-variable directly determines the behavior of the test. In the opposite direction, given an e-variable, we can associate with it a test obtained by thresholding. See Lemmas 4.5 and 4.7 and equations (4.3) and (4.6) for the details of these relationships. A second result worth mentioning is a Rao-Blackwell type-argument (Rao, 1945; Blackwell, 1947), whereby we demonstrate how to (weakly) improve any given test by conditioning on a sufficient statistic. This can be found in Section 5.2.
@@
Finally, we study two specific families of maps $\Gamma$ and give necessary and sufficient conditions on tests (and e-variables) to be $\Gamma$-admissible. In particular, we study both $\Gamma=\mathbf{U}$ and $\Gamma=\mathbf{C}$ where $\mathbf{U}$ is the set of all mappings from the data to the losses, and $\mathbf{C}$ is the set of constant mappings. For a point null and alternative, we give a full characterization of all $\mathbf{U}$-admissible tests in terms of e-variables for both randomized and binary tests—see Theorem 5.1 and Theorem 5.4. For $\mathbf{C}$, we give several necessary conditions for a randomized test to be admissible, and provide a complete characterization in the binary case—see Theorem 5.15. This final result recovers the Neyman-Pearson lemma (Neyman and Pearson, 1928) when instantiated with a single loss function. The curious reader can jump ahead to Table 2 for a diagramatic overview of our admissibility results.
@@ 1.2 Related Work
The post-hoc framework studied here is not the only proposal for how to resolve the issues with traditional hypothesis testing. Bayesian decision theory (Bernardo et al., 1994; Gelman et al., 1995) has a devoted following, where one replaces p-values and significance levels with posterior probabilities and Bayes factors, and makes decisions by minimizing posterior loss. A great benefit of the Bayesian paradigm is that it handles data-dependent loss functions for free, which we show formally in Appendix C where we explore the connection between our setting and the Bayesian one. The hefty price to pay is that one cannot get prior-free Type-I error probability guarantees.
@@
Besides Bayesian methods, other approaches to navigate around the drawbacks of traditional hypothesis testing involve changing what researchers report, e.g., switching p-values with confidence intervals (Gardner and Altman, 1986) or likelihood ratios (Royall, 2017). Post-hoc hypothesis testing has more in common with some of these than with others (it plays nicely with the law of likelihood for instance, since the likelihood ratio is an e-value), but importantly it is the only approach to seek a strict generalization of the traditional Neyman-Pearson paradigm. Given that most hypothesis testing in practice still relies on this paradigm, we believe it worthwhile to study how to expand the scope of its tools instead of casting them aside.
@@
Our work can be viewed as part of a more general focus on e-variables/e-values, which are simple yet surprisingly fruitful mathematical objects that have proven useful across a wide variety of statistical problems. While some authors had previously used them implicitly, it was not until recently (roughly 2020) that excitement grew concerning their applications and they were given a name. Since then, there has been growing interest in both developing the underlying theory of e-variables and deploying them in various problems. We refer to Grünwald et al. (2024), Ramdas and Wang (2025), and Vovk and Wang (2021) for introductions to e-values and their use in hypothesis testing.
@@
That e-values can play nicely with data-dependent parameters, thus overcoming some of the drawbacks of p-values, has been noted previously by several authors. The first work in this vein came from the multiple testing literature and worked with e-values only implicitly (Katsevich and Ramdas, 2020). As far as we are aware, the first result to explicitly link link e-values and data-dependent thresholds is Lemma 1 of Wang and Ramdas (2022) in the context of false discovery rate control. Then Grünwald (2024) put these connection on firm decision-theoretic ground and launched the original investigation into post-hoc hypothesis testing. In contradistinction to this paper, Grünwald considers multiple actions, whereas we consider only accept/reject decisions. Working under a different notion of admissibility, he gives a partial classification of all admissible tests in terms of e-variables. We further discuss his results and his notion of admissibility in Appendix B.
@@
As part of this work, Grünwald (2024) notes that e-value based confidence intervals and confidence distributions (using the so-called e-posterior, Grünwald, 2023) are safe under data-dependent selections of $\alpha$. Since then, a number of authors have noted the promise of e-values in enabling testing and estimating when significance levels are data-dependent (Xu et al., 2024; Koning, 2023; Hemerik and Koning, 2024; Xu et al., 2025; Gauthier et al., 2025). In particular, Koning (2023) also studies post-hoc hypothesis testing and defines post-hoc p-values. He further relates post-hoc p-values to e-values and shows that any post-hoc test defines an e-value. An overview of post-hoc testing and decision-making with e-values was recently given in Ramdas and Wang (2025, Chapter 4).
@@
Besides Grünwald (2024), several other authors have studied questions concerning the admissibility of e-variables. These include Wang (2025), Vovk and Wang (2020), and Ramdas et al. (2020). However, their notions of admissibility either differ from ours or the settings are incomparable. Wang (2025) and Vovk and Wang (2020) are concerned with designing e-values that are as large as possible under the alternative, and Ramdas et al. (2020) are concerned with admissible anytime-valid inference.
@@ 2 Post-hoc hypothesis testing
We begin with a standard hypothesis testing setup. Let $\mathcal{P}$ and $\mathcal{Q}$ be two disjoint families of probability measures on some measurable space $(\Omega,\mathcal{F})$. We observe data encoded by a random variable $X:\Omega\to\mathcal{X}$ taking values in some topological space $\mathcal{X}$ whose behavior is governed by some unknown distribution $R\in\mathcal{P}\cup\mathcal{Q}$. We are testing $\mathcal{P}$ against $\mathcal{Q}$, meaning that we are tasked with deciding between the null hypothesis $H_{0}:R\in\mathcal{P}$ and the alternative $H_{1}:R\in\mathcal{Q}$. We either reject or sustain the null hypothesis $H_{0}$. If the former, we are implicitly adopting the alternative.
@@
In order to formalize the notion of post-hoc significance levels, we follow Wald (1939) and Grünwald (2024) and frame hypothesis testing in terms of expected losses instead of error probabilities. For an index set $\mathcal{B}\subseteq\mathbb{R}_{\geq 0}$ which we call a set of scenarios, let $\{L_{b}\}_{b\in\mathcal{B}}$ be a set of loss functions, where $L_{b}:\{0,1\}\times\{0,1\}\to\mathbb{R}_{>0}$. The first argument denotes whether we are considering the loss under the null or alternative: $L_{b}(0,\cdot)$ is loss under the former and $L_{b}(1,\cdot)$ that under the latter. The second argument denotes whether we sustain (0) or reject (1) the null. We will assume that $L_{b}(0,0)=L_{b}(1,1)=0$ for each $b\in\mathcal{B}$. In other words, making the correct decision (sustaining when $R\in\mathcal{P}$ and rejecting when $R\in\mathcal{Q}$) incurs zero loss.
@@
Next we define the analogue of a hypothesis test in this post-hoc setting. Our decision to sustain or reject the null takes the form of a test family, which sees both the data and the loss, and then makes a (possibly randomized) decision. We emphasize that while we will allow the loss to be data-dependent, the test knows which loss function was chosen when making a decision. If it were otherwise, a test would have to be needlessly conservative. This is also the case in the traditional setting: the test knows the parameter $\alpha$.
@@
When studying binary test families, we will assume that $\mathcal{P}$ and $\mathcal{Q}$ consist of continuous distributions. Even in the traditional setting, we know that hypothesis tests must generally be randomized for discrete distributions so it makes little sense to study binary test families in this case.
@@
Risk is the equivalent of type-I error probability in the standard Neyman-Pearson paradigm. Indeed, consider the case of a single loss $\mathcal{B}=\{b_{0}\}$. Then $\mathcal{R}_{\mathcal{P}}(\delta)=\sup_{P\in\mathcal{P}}\mathbb{E}_{P}L_{b_{0}}(0,\delta(X,{b_{0}}))=L_{b_{0}}(0,1)\sup_{P\in\mathcal{P}}\mathbb{E}_{P}[\delta(X,b_{0})]$. A risk bounded by 1 is thus identical to a type-I error probability bounded by $\alpha>0$ if we take $L_{b_{0}}(0,1)=1/\alpha$. The upper bound of 1 on the risk is arbitrary. It can be replaced by any positive constant and our results will scale accordingly. See for instance Ramdas and Wang (2025, Section 4) which replaces 1 with a constant $\beta>0$.
While risk recovers type-I error probability when there is a single loss, risk is of course a different metric than an error probability in general. A framework built on replacing error probabilities with risk may therefore make some readers uncomfortable. We provide a longer defense of risk in Appendix D, but let us make two points here. First, risk is actually the norm in decision theory. Second, it is unclear how else to formulate a notion of post-hoc hypothesis testing.

### v3

@@ Abstract
The validity of classical hypothesis testing requires the significance level $\alpha$ be fixed before any statistical analysis takes place. This is a stringent requirement. For instance, it prohibits updating $\alpha$ during (or after) an experiment due to changing concern about the cost of false positives, or to reflect unexpectedly strong evidence against the null. Perhaps most disturbingly, witnessing a p-value $p\ll\alpha$ vs $p=\alpha-\epsilon$ for tiny $\epsilon>0$ has no (statistical) relevance for any downstream decision-making. Following recent work of grunwald2024beyond, we develop a theory of post-hoc hypothesis testing, enabling $\alpha$ to be chosen after seeing and analyzing the data. To study “good” post-hoc tests we introduce $\Gamma$-admissibility, where $\Gamma$ is a set of adversaries which map the data to a significance level. We classify the set of $\Gamma$-admissible rules for various sets $\Gamma$, showing they must be based on e-values, and recover the Neyman-Pearson lemma when $\Gamma$ is the constant map.
@@ 1 Introduction
In other words, the current paradigm of hypothesis testing requires that the significance level $\alpha$ be chosen independently of (thus without looking at) the data. This is well-known to statisticians, but continues to be a source of frustration and confusion for practitioners, and the epidemiologist’s mistake above is unfortunately all too common (bakan1966test; dar1994misuse; gigerenzer2004mindless). This fact is limiting: once the data has been gathered, if the selected value of $\alpha$ gives a vacuous or uninteresting result, nothing more can be done. The same data should not be used in any further analysis of the same question (or, technically, even a separate question motivated by the first [^2] This—admittedly philosophically thorny and somewhat controversial—reliance of p-values on counterfactuals is one of their many restrictive features (pratt1964foundations; wagenmakers2007practical).).
The difficulties do not end there. Requiring that $\alpha$ be chosen independently of the data introduces a fundamental tension between evidence and decision-making. Namely, upon observing a p-value $p\leq\alpha$, it is irrelevant whether $p$ is roughly the same as $\alpha$ or significantly smaller. For example, if $\alpha=0.05$, then observing $p=0.04$ and $p=10^{-6}$ have the same decision-theoretic consequences. As pointed out recently by grunwald2024beyond, this is a frustrating fact: the p-value is ostensibly a measure of evidence against the null hypothesis, yet any $p$-value below $\alpha$ has no formal relevance for any downstream decision task.
This limitation serves as an uncomfortable reminder that current statistical practice consists of an odd combination of both Ronald Fisher’s perspective and that of Jerzy Neyman and Egon Pearson. Whereas Fisher proposed the p-value as measure of evidence (fisher1925theory; fisher1935logic), Neyman and Pearson were focused on decision making (neyman1928use). The fact that we both report the p-value and make a decision at a prespecified significance level is an awkward methodological hybrid that none of these pioneers promoted.
This hybrid approach reflects (or perhaps introduces) confusion over the proper interpretation of p-values, significance levels, and hypothesis testing in general. And such confusion has substantial practical consequences. “The problem of roving alphas” as goodman1993p puts it, i.e., choosing $\alpha$ as a function of the data, is well-documented in business and management science (hubbard2003p; hubbard2011widespread), medicine (goodman1999toward), and psychology (gigerenzer2014superego). What are statisticians to do in the face of this mistake?
@@
Faced with such examples, in this work we consider an alternative response to the problem of roving alphas. Instead of pointing out flaws in the epidemiologist’s approach, we focus on building new statistical tools which function both as measures of evidence and allow for mathematically sound downstream decision-making. In particular, we develop the theory of post-hoc hypothesis testing (grunwald2024beyond), which allows for data-dependent significance levels. The catch is that the guarantees that an analyst can give on their procedure changes under this theory, moving from error probabilities to expected losses. If $\alpha$ is fixed beforehand, however, then the theory of post-hoc testing recovers the standard theory of hypothesis testing. Thus, in that sense, post-hoc hypothesis testing is a strict generation of the classical framework. Our focus in this paper is to define and classify “admissible test families”, which are the analogue of uniformly most powerful tests in the post-hoc setting.
One can situate our work as belonging to a recent push to resolve several issues with traditional statistical tools. The inability to handle post-hoc significance levels is just one problem with much of modern statistics—others include the inability to handle optional stopping and optional continuation. The burgeoning area of sequential, anytime-valid inference (SAVI) (ramdas2023game) is focused on such issues. A fundamental tool in this line of work is the e-value (grunwald2024safe; ramdas2024hypothesis), an object which will also play a crucial role in this paper. Overall, we view our work here as an extension of SAVI: broadly, as an attempt to refine existing statistical technology to make life easier for practitioners and to mitigate the risk of statistical malpractice.
@@ 1.1 Contributions
Following grunwald2024safe, in order to formulate a theory of post-hoc hypothesis testing we begin by following wald1939contributions and recast the usual notions of type-I and type-II error probabilities in terms of expectations over data-dependent type-I and type-II loss functions. We require that a test has bounded type-I loss in expectation, where the type-I loss may be chosen adversarially. Such a test is called type-I risk safe, a notion which generalizes the property of a test having a bounded type-I error probability.
At this point we depart from the setting studied by grunwald2024beyond and introduce a different notion of admissibility. We summarize our contributions below. While the general setting and several of the results are presented for composite hypotheses, we are mainly focused on testing point nulls and alternatives.
Throughout the paper it will often be useful to remark on the differences or similarities between our post-hoc setting and the traditional, non-post-hoc setting. Henceforth, we will often refer to the latter as the “traditional setting” or sometimes as the “Neyman-Pearson paradigm.” We refer to Neyman and Pearson because it was they who introduced type-I and type-II errors and the notion of a uniformly most powerful test (neyman1928use; neyman1933ix), concepts that we build upon and generalize in this work.
@@
Definition 2.4 introduces the notion of $\Gamma$-admissibility, where $\Gamma$ is a set of functions which map the data to a type-II loss function. This generalizes the definition of a uniformly most powerful test to the post-hoc setting (Observation 2.6 makes this precise). Informally, a test is $\Gamma$-admissible if there is no test that has smaller expected type-II loss under each distribution in the alternative and across all mappings in $\Gamma$. Our definition of $\Gamma$-admissibility departs from the notion of admissibility studied by grunwald2024beyond, which requires that a test be dominated with probability 1 under each distribution in the alternative to be inadmissible. We require that a test be dominated only in expectation, making fewer tests admissible.
@@
We give an extension of the likelihood ratio test—shown to be uniformly most powerful in the traditional paradigm (neyman1933ix)—to a post-hoc setting and prove that the extension is $\Gamma$-admissible for any $\Gamma$. This is in Section 3. This shows that our definition of admissibility is not too strict; the set of $\Gamma$-admissible tests is non-empty. This investigation also sheds light on an interesting phenomenon in the post-hoc setting: Even for continuous distributions, if randomization is allowed then admissible tests will always be randomized. More specifically, there will be losses on which the test is randomized. Since randomized tests may not be kosher for all applications, this motivates explicitly defining and studying the admissibility of binary post-hoc tests which reject or accept with probability one.
@@
Moving beyond the likelihood ratio test, we prove several general properties for all $\Gamma$-admissible tests for any $\Gamma$. We study both randomized and binary tests. Of particular note, we associate with each test an e-variable and show how any admissible test can be written as a function of the associated e-variable. Thus, the e-variable directly determines the behavior of the test. In the opposite direction, given an e-variable, we can associate with it a test obtained by thresholding. See Lemmas 4.5 and 4.7 and equations (4.3) and (4.6) for the details of these relationships. A second result worth mentioning is a Rao-Blackwell type-argument (radhakrishna1945information; blackwell1947conditional), whereby we demonstrate how to (weakly) improve any given test by conditioning on a sufficient statistic. This can be found in Section 5.2.
@@
Finally, we study two specific families of maps $\Gamma$ and give necessary and sufficient conditions on tests (and e-variables) to be $\Gamma$-admissible. In particular, we study both $\Gamma=\mathbf{U}$ and $\Gamma=\mathbf{C}$ where $\mathbf{U}$ is the set of all mappings from the data to the losses, and $\mathbf{C}$ is the set of constant mappings. For a point null and alternative, we give a full characterization of all $\mathbf{U}$-admissible tests in terms of e-variables for both randomized and binary tests—see Theorem 5.1 and Theorem 5.4. For $\mathbf{C}$, we give several necessary conditions for a randomized test to be admissible, and provide a complete characterization in the binary case—see Theorem 5.13. This final result recovers the Neyman-Pearson lemma (neyman1928use) when instantiated with a single loss function. The curious reader can jump ahead to Table 2 for a diagramatic overview of our admissibility results.
@@ 1.2 Related Work
The post-hoc framework studied here is not the only proposal for how to resolve the issues with traditional hypothesis testing. Bayesian decision theory (bernardo1994bayesian; gelman1995bayesian) has a devoted following, where one replaces p-values and significance levels with posterior probabilities and Bayes factors, and makes decisions by minimizing posterior loss. A great benefit of the Bayesian paradigm is that it handles data-dependent loss functions for free, which we show formally in Appendix C where we explore the connection between our setting and the Bayesian one. The hefty price to pay is that one cannot get prior-free Type-I error probability guarantees.
Another avenue of related work is the subfield of imprecise probability. Here one tries to avoid making claims with false precision by demanding that decisions be defensible over a set of models and/or priors (as done recently in, e.g., chau2025credal for two-sample testing). E-values may already be considered as a tool of imprecise probability, as they are closed under convex combinations of the null (sometimes called “credal sets”). Moreover, the e-posterior (grunwald2023posterior) was recently connected to credal sets (caprio2025joys, Proposition 10). While we do not pursue further connections here, we believe this to be an interesting direction.
@@
Besides imprecise probability and Bayesian methods, other approaches to navigate around the drawbacks of traditional hypothesis testing involve changing what researchers report, e.g., switching p-values with confidence intervals (gardner1986confidence) or likelihood ratios (royall2017statistical). Post-hoc hypothesis testing has more in common with some of these than with others (it plays nicely with the law of likelihood for instance, since the likelihood ratio is an e-value), but importantly it is the only approach to seek a strict generalization of the traditional Neyman-Pearson paradigm. Given that most hypothesis testing in practice still relies on this paradigm, we believe it worthwhile to study how to expand the scope of its tools instead of casting them aside.
@@
Our work can be viewed as part of a more general focus on e-variables/e-values, which are simple yet surprisingly fruitful mathematical objects that have proven useful across a wide variety of statistical problems. While some authors had previously used them implicitly, it was not until recently (roughly 2020) that excitement grew concerning their applications and they were given a name. Since then, there has been growing interest in both developing the underlying theory of e-variables and deploying them in various problems. We refer to grunwald2024safe, ramdas2024hypothesis, and vovk2021values for introductions to e-values and their use in hypothesis testing.
@@
That e-values can play nicely with data-dependent parameters, thus overcoming some of the drawbacks of p-values, has been noted previously by several authors. The first work in this vein came from the multiple testing literature and worked with e-values only implicitly (katsevich2020simultaneous). As far as we are aware, the first result to explicitly link link e-values and data-dependent thresholds is Lemma 1 of wang2022false in the context of false discovery rate control. Then grunwald2024beyond put these connection on firm decision-theoretic ground and launched the original investigation into post-hoc hypothesis testing. In contradistinction to this paper, Grünwald considers multiple actions, whereas we consider only accept/reject decisions. Working under a different notion of admissibility, he gives a partial classification of all admissible tests in terms of e-variables. We further discuss his results and his notion of admissibility in Appendix B.
@@
As part of this work, grunwald2024beyond notes that e-value based confidence intervals and confidence distributions (using the so-called e-posterior (grunwald2023posterior)) are safe under data-dependent selections of $\alpha$. Since then, a number of authors have noted the promise of e-values in enabling testing and estimating when significance levels are data-dependent (xu2024post; koning2023markov; hemerik2024choosing; xu2025bringing; gauthier2025values). In particular, koning2023markov also studies post-hoc hypothesis testing and defines post-hoc p-values. He further relates post-hoc p-values to e-values and shows that any post-hoc test defines an e-value. An overview of post-hoc testing and decision-making with e-values was recently given in ramdas2024hypothesis.
@@
Besides grunwald2024beyond, several other authors have studied questions concerning the admissibility of e-variables. These include wang2024only, vovk2020combining, and ramdas2020admissible. However, their notions of admissibility either differ from ours or the settings are incomparable. wang2024only and vovk2020combining are concerned with designing e-values that are as large as possible under the alternative, and ramdas2020admissible are concerned with admissible anytime-valid inference.
@@ 2 Post-hoc hypothesis testing
We begin with a standard hypothesis testing setup. Let $\mathcal{P}$ and $\mathcal{Q}$ be two disjoint families of probability measures on some measurable space $(\Omega,\mathcal{F})$. We observe data encoded by a random variable $X:\Omega\to\mathcal{X}$ taking values in some space $\mathcal{X}$ whose behavior is governed by some unknown distribution $R\in\mathcal{P}\cup\mathcal{Q}$. We are testing $\mathcal{P}$ against $\mathcal{Q}$, meaning that we are tasked with deciding between the null hypothesis $H_{0}:R\in\mathcal{P}$ and the alternative $H_{1}:R\in\mathcal{Q}$. We either reject or sustain the null hypothesis $H_{0}$. If the former, we are implicitly adopting the alternative.
@@
In order to formalize the notion of post-hoc significance levels, we follow wald1939contributions and grunwald2024beyond and frame hypothesis testing in terms of expected losses instead of error probabilities. For an index set $\mathcal{B}\subseteq\mathbb{R}_{\geq 0}$ which we call a set of scenarios, let $\{L_{b}\}_{b\in\mathcal{B}}$ be a set of loss functions, where $L_{b}:\{0,1\}\times\{0,1\}\to\mathbb{R}_{>0}$. The first argument denotes whether we are considering the loss under the null or alternative: $L_{b}(0,\cdot)$ is loss under the former and $L_{b}(1,\cdot)$ that under the latter. The second argument denotes whether we sustain (0) or reject (1) the null. We will assume that $L_{b}(0,0)=L_{b}(1,1)=0$ for each $b\in\mathcal{B}$. In other words, making the correct decision (sustaining when $R\in\mathcal{P}$ and rejecting when $R\in\mathcal{Q}$) incurs zero loss.
@@
Next we define the analogue of a hypothesis test in this post-hoc setting. Our decision to sustain or reject the null takes the form of a test family, which sees both the data and the loss, and then makes a (possibly randomized) decision. We emphasize that while we will allow the loss to be data-dependent, the test takes this loss as an argument. If it were otherwise, a test would have to be needlessly conservative. This is also the case in the traditional setting: the test knows the parameter $\alpha$.
@@
When studying binary test families, we will assume that $\mathcal{P}$ and $\mathcal{Q}$ consist of continuous distributions. Even in the traditional setting, we know that powerful hypothesis tests must generally be randomized for discrete distributions. Further, in the case of discrete distributions and binary tests, there is no neat classification of UMP tests as far as we are aware. Thus, studying admissibility in this case becomes quite challenging.
@@
Risk is the equivalent of type-I error probability in the standard Neyman-Pearson paradigm. Indeed, consider the case of a single loss $\mathcal{B}=\{b_{0}\}$. Then $\mathcal{R}_{\mathcal{P}}(\delta)=\sup_{P\in\mathcal{P}}\mathbb{E}_{P}L_{b_{0}}(0,\delta(X,{b_{0}}))=L_{b_{0}}(0,1)\sup_{P\in\mathcal{P}}\mathbb{E}_{P}[\delta(X,b_{0})]$. A risk bounded by 1 is thus identical to a type-I error probability bounded by $\alpha>0$ if we take $L_{b_{0}}(0,1)=1/\alpha$. The upper bound of 1 on the risk is arbitrary. It can be replaced by any positive constant and our results will scale accordingly. See for instance ramdas2024hypothesis which replaces 1 with a constant $\beta>0$.
While risk recovers type-I error probability when there is a single loss, risk is of course a different metric than an error probability in general. A framework built on replacing error probabilities with risk may therefore make some readers uncomfortable. We provide a longer defense of risk in Appendix D, but let us make two points here. First, risk is actually the norm in decision theory. Second, type-I risk seems to us the most straightforward generalization of type-I error control. [^3] For the interested reader, koning2023markov explores two other options, so-called “geometric” and “arithmetic” post-hoc validity.